If cosx+2cos(x+240°)=1

Determine the general equations


1
Expert's answer
2021-05-07T09:15:38-0400


Cos(x) + cos (x+240o) =1

"\implies"

"cos(x)+cos(x+{24\pi \over18}) = 1"


"cos(x)+cos(x+{24\pi \over18}) = cos(0)"



We know that "\boxed{cos(a+b) = cos(a)cos(b)-sin(a)sin(b)}"

"\implies"

"\boxed{cos(x)+cos(x)cos(24\pi/18)-sin(x)sin(24\pi/18)=cos(0)}"

On solving this

We got

"\boxed{x = \frac{1}{3}(6\pi n +\pi)}"